OpenStudy (anonymous):

Solve using systems of equations and without using systems of equations. Larry and Louise are siblings. Louise has twice as many brothers as she does sisters. Larry has the same number of brothers as sisters. How many boys and how many girls are in the family?

OpenStudy (anonymous):

System: (x for boys, y for girls) y=(1/2)x+1 and x=y+1 x=y+1 -1 x-1=y y=(1/2)x+1 x-1=(1/2)x+1 -1 x-2=(1/2)x times 2 2x-4=x -2x -4=x-2x -4=-x times -1 4=x x=y+1 4=y+1 -1 3=y (Sorry, I don't know how to go about this without systems.)

OpenStudy (anonymous):

thank you